A tank contains 200 liters of fresh water .Brine containing 2.5 N/liter of dissolved salt runs into the tank at the rate of 8 liters/min and the mixture kept uniform by stirring runs out at 4 liters per minute .Find the amount when the tank contains 240 liters of brine. The concentration of the salt in the tank after 25 minutes amounts to how much


1
Expert's answer
2021-11-09T15:11:13-0500

Solution;

Let A(t) be the amount of salt in the tank at any time t.

"\frac{dA}{dt}=R_{in}-R_{out}" .....(i)

From the given information;

"R_{in}=2.5\u00d78=20N\/min"

"R_{out}=\frac{4×A}{200+4t}"

"\frac{dA}{dt}=20-\frac{4A}{200+4t}"

Rewrite as;

"\frac{dA}{dt}+\frac{A}{50+t}=20"

This is a first order differential equation;

The integrating factor:

"I.F=e^{\frac{1}{50+t}dt}" ="e^{log(50+t)}=50+t"

Solution of the equation is;

"A×I.F=\int (I.F×20)dt +c"

"A(50+t)=\int20(50+t)dt+c"

"A(50+t)=20(50t+\frac{t^2}{2})+c"

"A(50+t)=1000t+10t^2+c"

Initial condition;

A(0)=0

Apply;

"0(50+t)=0+0+c"

Hence ; c=0

And;

"A=\frac{1000t+10t^2}{50+t}" ....(2)

The tank contains 240ltrs of brine after;

"\frac{240-200}{4}=10 minutes"

After 10 minutes;

"A(10)=\frac{(1000×10)+(10×10^2)}{50+10}"

"A(10)=188.83N"

Now concentration of salt in the tank after 25 minutes;

"A(25)=\frac{(1000×25)-(10×25^2)}{50+25}"

"A(25)=416.66N"